how do you solve the equation for 54x3-15x2-6x=0 I'm working on polynomials and this question has me lost.
Answer by nerdybill(7384)

54x^3-15x2-6x=0
begin by factoring out what is common among all three terms:
3x(18x^2-5x-2)=0
Now, factor the polynomial inside the parentheses:
3x(18x^2-5x-2)=0
rewrite the middle term as:
3x(18x^2-9x+4x-2)=0
group
3x((18x^2-9x)+(4x-2))=0
3x(9x(2x-1)+2(2x-1))=0
3x((2x-1)(9x+2))=0
3x(2x-1)(9x+2)=0
x = {0, 1/2, -2/9}